3. (20 points) Find the solution to the differential equation y sin(y) dx + x(sin y - ycos y) dy = 0
Differential Equations
Answer #1

y\sin(y)\,dx+x(\sin(y)-y\cos(y))\,dy=0

Is M\,dx+N\,dy=0 with/ M=y\sin(y),N=x(\sin(y)-y\cos(y))

So M_y=y\cos(y)+\sin(y),N_x=\sin(y)-y\cos(y)

We have \frac{M_y-N_x}{-M}=\frac{y\cos(y)+\sin(y)-\sin(y)+y\cos(y)}{-y\sin(y)}=\frac{2y\cos(y)}{-y\sin(y)}=-2\cot(y)

Integrating factor is \exp\left(\int (-2\cot(y)) \right )=\exp\left(-2\ln(\sin(y)) \right )=\csc^2(y)

Multiplying by thi/s factor we get

y\csc(y)\,dx+x(\csc(y)-y\csc(y)\cot(y))\,dy=0

We can check that this equation is now exact as \frac{\partial }{\partial y}\left(y\csc(y) \right )=\frac{\partial }{\partial x}\left(x(\csc(y)-y\csc(y)\cot(y)) \right )

And so the solution is u(x,y)=C such that u_x(x,y)=y\csc(y)\Rightarrow u(x,y)=xy\cos(y)+g(y)

Substituting in u_y(x,y)=x(\csc(y)-y\csc(y)\cot(y))\Rightarrow g'(y)=0

So our solution is simply {\color{Red} xy\csc(y)=C}

\blacksquare